# TimeBridge: Non-Stationarity Matters for Long-term Time Series Forecasting

### v2

@@ E.1 Main Experiments
ETTm1 96 0.297 0.353 0.334 0.368 0.280 0.335 0.320 0.357 0.293 0.346 0.316 0.373 0.379 0.419 0.292 0.346 0.316 0.362 0.338 0.375 0.299 0.343
192 0.333 0.375 0.377 0.391 0.317 0.359 0.361 0.381 0.333 0.370 0.377 0.411 0.426 0.441 0.332 0.368 0.363 0.390 0.374 0.387 0.335 0.365
336 0.366 0.399 0.426 0.420 0.354 0.382 0.390 0.404 0.369 0.392 0.431 0.442 0.445 0.459 0.365 0.391 0.408 0.426 0.410 0.411 0.369 0.386
720 0.414 0.423 0.491 0.459 0.405 0.413 0.454 0.441 0.416 0.420 0.600 0.547 0.543 0.490 0.416 0.417 0.481 0.476 0.478 0.450 0.425 0.421
Avg. 0.353 0.388 0.407 0.410 0.339 0.372 0.381 0.395 0.353 0.382 0.431 0.443 0.448 0.452 0.351 0.381 0.392 0.414 0.400 0.406 0.357 0.379
ETTm2 96 0.158 0.249 0.180 0.264 0.162 0.253 0.175 0.258 0.166 0.256 0.275 0.358 0.203 0.287 0.166 0.256 0.203 0.287 0.187 0.267 0.167 0.260
192 0.215 0.291 0.250 0.309 0.219 0.291 0.237 0.299 0.223 0.296 0.345 0.400 0.269 0.328 0.222 0.293 0.262 0.326 0.249 0.309 0.224 0.303
336 0.263 0.323 0.311 0.348 0.270 0.326 0.298 0.340 0.274 0.329 0.657 0.528 0.325 0.366 0.272 0.324 0.305 0.353 0.321 0.351 0.281 0.342
720 0.348 0.376 0.412 0.407 0.358 0.380 0.391 0.396 0.362 0.385 1.208 0.753 0.421 0.415 0.351 0.381 0.389 0.407 0.408 0.403 0.397 0.421
Avg. 0.246 0.310 0.288 0.332 0.252 0.313 0.275 0.323 0.256 0.317 0.621 0.510 0.305 0.349 0.253 0.314 0.290 0.343 0.291 0.333 0.267 0.332
ETTh1 96 0.358 0.392 0.386 0.405 0.357 0.388 0.375 0.400 0.370 0.400 0.405 0.426 0.376 0.419 0.368 0.394 0.398 0.427 0.384 0.402 0.375 0.399
192 0.388 0.411 0.441 0.436 0.397 0.412 0.429 0.421 0.413 0.429 0.419 0.444 0.420 0.448 0.405 0.413 0.430 0.453 0.436 0.429 0.405 0.416
336 0.401 0.419 0.487 0.458 0.409 0.422 0.484 0.458 0.422 0.440 0.440 0.461 0.459 0.465 0.391 0.412 0.440 0.460 0.491 0.469 0.439 0.443
720 0.447 0.458 0.503 0.491 0.432 0.455 0.498 0.482 0.447 0.468 0.519 0.524 0.506 0.507 0.450 0.461 0.491 0.509 0.521 0.500 0.472 0.490
Avg. 0.399 0.420 0.454 0.447 0.399 0.419 0.447 0.440 0.413 0.434 0.446 0.464 0.440 0.460 0.404 0.420 0.440 0.462 0.458 0.450 0.423 0.437
ETTh2 96 0.295 0.354 0.297 0.349 0.272 0.333 0.289 0.341 0.274 0.337 0.628 0.563 0.346 0.388 0.263 0.332 0.332 0.377 0.340 0.374 0.289 0.353
192 0.351 0.389 0.380 0.400 0.335 0.375 0.372 0.392 0.314 0.382 0.703 0.624 0.429 0.439 0.320 0.374 0.422 0.441 0.402 0.414 0.383 0.418
336 0.351 0.397 0.428 0.432 0.325 0.377 0.386 0.414 0.329 0.384 0.827 0.675 0.496 0.487 0.313 0.376 0.447 0.474 0.452 0.452 0.448 0.465
720 0.388 0.436 0.427 0.445 0.375 0.417 0.412 0.434 0.379 0.422 1.181 0.840 0.463 0.474 0.392 0.433 0.442 0.467 0.462 0.468 0.605 0.551
Avg. 0.346 0.394 0.383 0.407 0.327 0.376 0.364 0.395 0.324 0.381 0.835 0.675 0.434 0.447 0.322 0.379 0.411 0.440 0.414 0.427 0.431 0.447
Weather 96 0.143 0.192 0.174 0.214 0.147 0.194 0.163 0.209 0.149 0.198 0.153 0.217 0.217 0.296 0.149 0.200 0.161 0.229 0.172 0.220 0.176 0.237
192 0.185 0.235 0.221 0.254 0.192 0.239 0.208 0.250 0.194 0.241 0.197 0.269 0.276 0.336 0.196 0.245 0.220 0.281 0.219 0.261 0.220 0.282
336 0.237 0.277 0.278 0.296 0.244 0.279 0.251 0.287 0.245 0.282 0.495 0.515 0.339 0.380 0.238 0.277 0.278 0.331 0.280 0.306 0.265 0.319
720 0.307 0.330 0.358 0.349 0.318 0.330 0.339 0.341 0.314 0.334 0.526 0.542 0.403 0.428 0.314 0.334 0.311 0.356 0.365 0.359 0.323 0.362
Avg. 0.218 0.259 0.258 0.279 0.225 0.261 0.240 0.271 0.226 0.264 0.343 0.386 0.309 0.360 0.224 0.264 0.243 0.299 0.259 0.287 0.246 0.300
Electricity 96 0.118 0.218 0.148 0.240 0.127 0.219 0.153 0.247 0.129 0.222 0.187 0.283 0.183 0.297 0.129 0.226 0.164 0.269 0.168 0.272 0.140 0.237
192 0.142 0.237 0.162 0.253 0.145 0.237 0.166 0.256 0.147 0.240 0.258 0.330 0.195 0.308 0.143 0.239 0.177 0.285 0.184 0.289 0.153 0.249
336 0.156 0.252 0.178 0.269 0.162 0.255 0.185 0.277 0.163 0.259 0.323 0.369 0.212 0.313 0.161 0.259 0.193 0.304 0.198 0.300 0.169 0.267
720 0.179 0.278 0.225 0.317 0.200 0.290 0.225 0.310 0.197 0.290 0.404 0.423 0.231 0.343 0.191 0.286 0.212 0.321 0.220 0.320 0.203 0.301
Avg. 0.149 0.246 0.178 0.270 0.159 0.250 0.182 0.272 0.159 0.253 0.293 0.351 0.205 0.315 0.156 0.253 0.187 0.295 0.192 0.295 0.166 0.264
Traffic 96 0.340 0.240 0.395 0.268 0.351 0.238 0.462 0.285 0.360 0.249 0.512 0.290 0.562 0.349 0.368 0.253 0.519 0.309 0.593 0.321 0.410 0.282
192 0.343 0.250 0.417 0.276 0.374 0.248 0.473 0.296 0.379 0.256 0.523 0.297 0.562 0.346 0.379 0.261 0.537 0.315 0.617 0.336 0.423 0.287
336 0.363 0.257 0.433 0.283 0.386 0.253 0.498 0.296 0.392 0.264 0.530 0.300 0.570 0.323 0.397 0.270 0.534 0.313 0.629 0.336 0.436 0.296
720 0.393 0.271 0.467 0.302 0.421 0.278 0.506 0.313 0.432 0.286 0.573 0.313 0.596 0.368 0.440 0.296 0.577 0.325 0.640 0.350 0.466 0.315
Avg. 0.360 0.255 0.428 0.282 0.383 0.254 0.484 0.297 0.391 0.264 0.535 0.300 0.573 0.347 0.396 0.270 0.542 0.316 0.620 0.336 0.434 0.295
Solar 96 0.161 0.224 0.203 0.237 0.179 0.246 0.189 0.259 0.190 0.273 0.181 0.240 0.209 0.330 0.202 0.263 0.190 0.250 0.285 0.330 0.289 0.377
192 0.177 0.237 0.233 0.261 0.205 0.265 0.222 0.283 0.204 0.302 0.196 0.252 0.274 0.400 0.223 0.279 0.225 0.270 0.309 0.342 0.319 0.397
336 0.188 0.244 0.248 0.273 0.210 0.270 0.231 0.292 0.212 0.293 0.216 0.243 0.338 0.439 0.241 0.292 0.250 0.301 0.335 0.365 0.352 0.415
720 0.197 0.252 0.249 0.275 0.225 0.281 0.223 0.285 0.221 0.310 0.220 0.256 0.365 0.459 0.247 0.292 0.323 0.362 0.346 0.355 0.356 0.412
Avg. 0.181 0.239 0.233 0.262 0.205 0.265 0.216 0.280 0.207 0.294 0.204 0.248 0.296 0.407 0.228 0.282 0.247 0.296 0.319 0.348 0.329 0.400
Table 10: Full results of long-term forecasting from the original papers. All results are averaged across four different prediction lengths: $O\in\{96,192,336,720\}$. The best and second-best results are highlighted in bold and underlined, respectively.
Models TimeBridge iTransformer PDF TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1 96 0.297 0.353 0.300 0.353 0.277 0.337 0.291 0.340 0.293 0.346 0.310 0.361 0.326 0.390 0.292 0.346 0.314 0.360 0.338 0.375 0.299 0.343
192 0.333 0.375 0.345 0.382 0.316 0.364 0.327 0.365 0.333 0.370 0.363 0.402 0.365 0.415 0.332 0.368 0.359 0.387 0.371 0.387 0.335 0.365
336 0.366 0.399 0.374 0.398 0.346 0.381 0.360 0.381 0.369 0.392 0.408 0.430 0.392 0.425 0.365 0.391 0.398 0.413 0.410 0.411 0.369 0.386
720 0.414 0.423 0.429 0.430 0.402 0.409 0.415 0.417 0.416 0.420 0.600 0.547 0.446 0.458 0.416 0.417 0.459 0.464 0.478 0.450 0.425 0.421
Avg. 0.353 0.388 0.362 0.391 0.335 0.373 0.348 0.375 0.353 0.382 0.420 0.435 0.382 0.422 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 96 0.158 0.249 0.175 0.266 0.159 0.251 0.164 0.254 0.166 0.256 0.263 0.359 0.180 0.271 0.166 0.256 0.178 0.273 0.187 0.267 0.167 0.260
192 0.215 0.291 0.242 0.312 0.217 0.292 0.223 0.295 0.223 0.296 0.345 0.400 0.252 0.318 0.222 0.293 0.245 0.316 0.249 0.309 0.224 0.303
336 0.263 0.323 0.282 0.340 0.266 0.325 0.279 0.330 0.274 0.329 0.469 0.496 0.324 0.364 0.272 0.324 0.295 0.350 0.321 0.351 0.281 0.342
720 0.348 0.376 0.378 0.398 0.345 0.375 0.359 0.383 0.362 0.385 0.996 0.750 0.410 0.420 0.351 0.381 0.389 0.406 0.497 0.403 0.397 0.421
Avg. 0.246 0.310 0.269 0.329 0.247 0.311 0.256 0.315 0.256 0.317 0.518 0.501 0.292 0.343 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 96 0.358 0.392 0.386 0.405 0.356 0.391 0.361 0.390 0.370 0.400 0.386 0.426 0.376 0.415 0.368 0.394 0.396 0.427 0.384 0.402 0.375 0.399
192 0.388 0.411 0.424 0.440 0.390 0.413 0.409 0.414 0.413 0.429 0.413 0.442 0.423 0.446 0.405 0.413 0.430 0.453 0.557 0.436 0.405 0.416
336 0.401 0.419 0.449 0.460 0.402 0.421 0.430 0.429 0.422 0.440 0.440 0.461 0.444 0.462 0.391 0.412 0.433 0.458 0.491 0.469 0.439 0.443
720 0.447 0.458 0.495 0.487 0.432 0.455 0.445 0.460 0.447 0.468 0.519 0.524 0.469 0.492 0.450 0.461 0.474 0.508 0.521 0.500 0.472 0.490
Avg. 0.399 0.420 0.439 0.448 0.395 0.420 0.411 0.423 0.413 0.434 0.440 0.463 0.428 0.454 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 96 0.295 0.354 0.297 0.348 0.270 0.332 0.271 0.330 0.274 0.337 0.611 0.557 0.332 0.374 0.263 0.332 0.289 0.357 0.340 0.374 0.289 0.353
192 0.351 0.389 0.371 0.403 0.334 0.375 0.317 0.402 0.314 0.382 0.703 0.624 0.407 0.446 0.320 0.374 0.409 0.438 0.402 0.414 0.383 0.418
336 0.351 0.397 0.404 0.428 0.324 0.379 0.332 0.396 0.329 0.384 0.827 0.675 0.400 0.447 0.313 0.376 0.417 0.452 0.452 0.452 0.448 0.465
720 0.388 0.436 0.424 0.444 0.375 0.417 0.342 0.408 0.379 0.422 1.094 0.775 0.412 0.469 0.392 0.433 0.426 0.473 0.462 0.468 0.605 0.551
Avg. 0.346 0.394 0.374 0.406 0.326 0.376 0.316 0.384 0.324 0.381 0.809 0.658 0.388 0.434 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 96 0.143 0.192 0.159 0.208 0.143 0.193 0.147 0.197 0.149 0.198 0.146 0.212 0.217 0.296 0.149 0.200 0.161 0.226 0.172 0.220 0.152 0.237
192 0.185 0.235 0.200 0.248 0.188 0.236 0.189 0.239 0.194 0.241 0.195 0.261 0.275 0.329 0.196 0.245 0.220 0.283 0.219 0.261 0.220 0.282
336 0.237 0.277 0.253 0.289 0.240 0.279 0.241 0.280 0.245 0.282 0.252 0.311 0.339 0.377 0.238 0.277 0.275 0.328 0.280 0.306 0.265 0.319
720 0.307 0.330 0.321 0.338 0.308 0.328 0.310 0.330 0.314 0.334 0.318 0.363 0.389 0.409 0.314 0.334 0.311 0.356 0.365 0.359 0.323 0.362
Avg. 0.218 0.259 0.233 0.271 0.220 0.259 0.222 0.262 0.226 0.264 0.228 0.287 0.305 0.287 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 96 0.118 0.218 0.138 0.237 0.126 0.220 0.129 0.224 0.129 0.222 0.135 0.237 0.183 0.297 0.129 0.226 0.159 0.267 0.168 0.272 0.140 0.237
192 0.142 0.237 0.157 0.256 0.145 0.237 0.140 0.220 0.147 0.240 0.160 0.262 0.195 0.308 0.143 0.239 0.168 0.279 0.184 0.289 0.152 0.249
336 0.156 0.252 0.167 0.264 0.159 0.255 0.161 0.255 0.163 0.259 0.182 0.282 0.212 0.313 0.161 0.259 0.196 0.308 0.198 0.300 0.169 0.267
720 0.179 0.278 0.194 0.286 0.194 0.287 0.194 0.287 0.197 0.290 0.246 0.337 0.231 0.343 0.191 0.286 0.203 0.312 0.220 0.320 0.203 0.301
Avg. 0.149 0.246 0.164 0.261 0.156 0.250 0.156 0.246 0.159 0.253 0.181 0.279 0.205 0.315 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 96 0.340 0.240 0.363 0.265 0.350 0.239 0.360 0.249 0.360 0.249 0.512 0.282 0.562 0.349 0.368 0.253 0.508 0.301 0.593 0.321 0.410 0.282
192 0.343 0.250 0.385 0.273 0.363 0.247 0.375 0.250 0.379 0.256 0.501 0.273 0.562 0.346 0.379 0.261 0.536 0.315 0.617 0.336 0.423 0.287
336 0.363 0.257 0.396 0.277 0.376 0.258 0.385 0.270 0.392 0.264 0.507 0.279 0.570 0.323 0.397 0.270 0.525 0.310 0.629 0.336 0.436 0.296
720 0.393 0.271 0.445 0.312 0.419 0.279 0.430 0.281 0.432 0.286 0.571 0.301 0.596 0.368 0.440 0.296 0.571 0.323 0.640 0.350 0.466 0.315
Avg. 0.360 0.255 0.397 0.282 0.377 0.256 0.387 0.262 0.391 0.264 0.523 0.284 0.573 0.347 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 96 0.161 0.224 0.188 0.242 0.179 0.246 0.167 0.220 0.178 0.229 0.166 0.230 0.201 0.304 0.202 0.263 0.188 0.252 0.219 0.314 0.216 0.287
192 0.177 0.237 0.193 0.258 0.205 0.265 0.187 0.249 0.189 0.246 0.186 0.237 0.237 0.337 0.223 0.279 0.215 0.280 0.231 0.322 0.244 0.305
336 0.188 0.244 0.195 0.259 0.210 0.270 0.200 0.258 0.198 0.249 0.203 0.243 0.254 0.362 0.241 0.292 0.222 0.267 0.246 0.337 0.263 0.319
720 0.197 0.252 0.223 0.281 0.225 0.281 0.215 0.250 0.209 0.256 0.210 0.256 0.280 0.397 0.247 0.292 0.226 0.264 0.280 0.363 0.264 0.324
Avg. 0.181 0.239 0.200 0.260 0.205 0.265 0.192 0.244 0.194 0.245 0.191 0.242 0.243 0.350 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309

### v3

@@ E.1 Main Experiments
ETTm1 96 0.284 0.337 0.300 0.353 0.291 0.347 0.293 0.345 0.293 0.346 0.310 0.361 0.326 0.390 0.292 0.346 0.314 0.360 0.338 0.375 0.299 0.343
192 0.317 0.367 0.345 0.382 0.325 0.372 0.335 0.372 0.333 0.370 0.363 0.402 0.365 0.415 0.332 0.368 0.359 0.387 0.371 0.387 0.335 0.365
336 0.361 0.394 0.374 0.398 0.359 0.390 0.368 0.386 0.369 0.392 0.408 0.430 0.392 0.425 0.365 0.391 0.398 0.413 0.410 0.411 0.369 0.386
720 0.413 0.418 0.429 0.430 0.418 0.423 0.426 0.417 0.416 0.420 0.600 0.547 0.446 0.458 0.416 0.417 0.459 0.464 0.478 0.450 0.425 0.421
Avg. 0.344 0.379 0.362 0.391 0.348 0.383 0.355 0.380 0.353 0.382 0.420 0.435 0.382 0.422 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 96 0.157 0.243 0.175 0.266 0.169 0.258 0.165 0.256 0.166 0.256 0.263 0.359 0.180 0.271 0.166 0.256 0.178 0.273 0.187 0.267 0.167 0.260
192 0.217 0.285 0.242 0.312 0.229 0.299 0.225 0.298 0.223 0.296 0.345 0.400 0.252 0.318 0.222 0.293 0.245 0.316 0.249 0.309 0.224 0.303
336 0.269 0.321 0.282 0.340 0.280 0.333 0.277 0.332 0.274 0.329 0.469 0.496 0.324 0.364 0.272 0.324 0.295 0.350 0.321 0.351 0.281 0.342
720 0.348 0.378 0.378 0.398 0.349 0.384 0.360 0.387 0.362 0.385 0.996 0.750 0.410 0.420 0.351 0.381 0.389 0.406 0.497 0.403 0.397 0.421
Avg. 0.246 0.310 0.269 0.329 0.257 0.319 0.257 0.318 0.256 0.317 0.518 0.501 0.292 0.343 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 96 0.350 0.389 0.386 0.405 0.369 0.396 0.372 0.401 0.370 0.400 0.386 0.426 0.376 0.415 0.368 0.394 0.396 0.427 0.384 0.402 0.375 0.399
192 0.388 0.414 0.424 0.440 0.410 0.417 0.413 0.430 0.413 0.429 0.413 0.442 0.423 0.446 0.405 0.413 0.430 0.453 0.557 0.436 0.405 0.416
336 0.408 0.430 0.449 0.460 0.391 0.414 0.438 0.450 0.422 0.440 0.440 0.461 0.444 0.462 0.391 0.412 0.433 0.458 0.491 0.469 0.439 0.443
720 0.443 0.463 0.495 0.487 0.447 0.464 0.486 0.484 0.447 0.468 0.519 0.524 0.469 0.492 0.450 0.461 0.474 0.508 0.521 0.500 0.472 0.490
Avg. 0.397 0.424 0.439 0.448 0.404 0.423 0.427 0.441 0.413 0.434 0.440 0.463 0.428 0.454 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 96 0.271 0.331 0.297 0.348 0.272 0.334 0.281 0.351 0.274 0.337 0.611 0.557 0.332 0.374 0.263 0.332 0.289 0.357 0.340 0.374 0.289 0.353
192 0.335 0.370 0.371 0.403 0.325 0.369 0.349 0.387 0.314 0.382 0.703 0.624 0.407 0.446 0.320 0.374 0.409 0.438 0.402 0.414 0.383 0.418
336 0.371 0.402 0.404 0.428 0.319 0.373 0.366 0.413 0.329 0.384 0.827 0.675 0.400 0.447 0.313 0.376 0.417 0.452 0.452 0.452 0.448 0.465
720 0.387 0.425 0.424 0.444 0.395 0.433 0.401 0.436 0.379 0.422 1.094 0.775 0.412 0.469 0.392 0.433 0.426 0.473 0.462 0.468 0.605 0.551
Avg. 0.341 0.382 0.374 0.406 0.328 0.377 0.349 0.397 0.324 0.381 0.809 0.658 0.388 0.434 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 96 0.144 0.184 0.159 0.208 0.146 0.198 0.147 0.198 0.149 0.198 0.146 0.212 0.217 0.296 0.149 0.200 0.161 0.226 0.172 0.220 0.152 0.237
192 0.186 0.225 0.200 0.248 0.191 0.239 0.192 0.243 0.194 0.241 0.195 0.261 0.275 0.329 0.196 0.245 0.220 0.283 0.219 0.261 0.220 0.282
336 0.237 0.267 0.253 0.289 0.241 0.280 0.247 0.284 0.245 0.282 0.252 0.311 0.339 0.377 0.238 0.277 0.275 0.328 0.280 0.306 0.265 0.319
720 0.307 0.320 0.321 0.338 0.310 0.331 0.318 0.330 0.314 0.334 0.318 0.363 0.389 0.409 0.314 0.334 0.311 0.356 0.365 0.359 0.323 0.362
Avg. 0.219 0.249 0.233 0.271 0.222 0.262 0.226 0.264 0.226 0.264 0.228 0.287 0.305 0.287 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 96 0.120 0.214 0.138 0.237 0.132 0.234 0.153 0.256 0.129 0.222 0.135 0.237 0.183 0.297 0.129 0.226 0.159 0.267 0.168 0.272 0.140 0.237
192 0.142 0.237 0.157 0.256 0.148 0.248 0.168 0.269 0.147 0.240 0.160 0.262 0.195 0.308 0.143 0.239 0.168 0.279 0.184 0.289 0.152 0.249
336 0.156 0.252 0.167 0.264 0.165 0.266 0.189 0.291 0.163 0.259 0.182 0.282 0.212 0.313 0.161 0.259 0.196 0.308 0.198 0.300 0.169 0.267
720 0.179 0.278 0.194 0.286 0.197 0.296 0.228 0.320 0.197 0.290 0.246 0.337 0.231 0.343 0.191 0.286 0.203 0.312 0.220 0.320 0.203 0.301
Avg. 0.149 0.245 0.164 0.261 0.161 0.261 0.185 0.284 0.159 0.253 0.181 0.279 0.205 0.315 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 96 0.340 0.240 0.363 0.265 0.355 0.261 0.369 0.257 0.360 0.249 0.512 0.282 0.562 0.349 0.368 0.253 0.508 0.301 0.593 0.321 0.410 0.282
192 0.343 0.250 0.385 0.273 0.380 0.271 0.400 0.272 0.379 0.256 0.501 0.273 0.562 0.346 0.379 0.261 0.536 0.315 0.617 0.336 0.423 0.287
336 0.363 0.257 0.396 0.277 0.393 0.281 0.407 0.272 0.392 0.264 0.507 0.279 0.570 0.323 0.397 0.270 0.525 0.310 0.629 0.336 0.436 0.296
720 0.393 0.271 0.445 0.312 0.434 0.300 0.461 0.316 0.432 0.286 0.571 0.301 0.596 0.368 0.440 0.296 0.571 0.323 0.640 0.350 0.466 0.315
Avg. 0.360 0.255 0.397 0.282 0.391 0.278 0.409 0.279 0.391 0.264 0.523 0.284 0.573 0.347 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 96 0.161 0.224 0.188 0.242 0.165 0.238 0.179 0.232 0.178 0.229 0.166 0.230 0.201 0.304 0.202 0.263 0.188 0.252 0.219 0.314 0.216 0.287
192 0.177 0.237 0.193 0.258 0.184 0.254 0.201 0.259 0.189 0.246 0.186 0.237 0.237 0.337 0.223 0.279 0.215 0.280 0.231 0.322 0.244 0.305
336 0.188 0.244 0.195 0.259 0.191 0.263 0.190 0.256 0.198 0.249 0.203 0.243 0.254 0.362 0.241 0.292 0.222 0.267 0.246 0.337 0.263 0.319
720 0.197 0.252 0.223 0.281 0.199 0.262 0.203 0.261 0.209 0.256 0.210 0.256 0.280 0.397 0.247 0.292 0.226 0.264 0.280 0.363 0.264 0.324
Avg. 0.181 0.239 0.200 0.260 0.185 0.254 0.193 0.252 0.194 0.245 0.191 0.242 0.243 0.350 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309
Climate 96 1.040 0.492 1.044 0.494 1.042 0.498 1.341 0.578 1.354 0.519 1.083 0.508 1.290 0.620 1.100 0.514 1.150 0.574 1.274 0.524 1.108 0.536
192 1.049 0.494 1.055 0.496 1.051 0.494 1.447 0.622 1.388 0.572 1.115 0.532 1.447 0.693 1.125 0.523 1.159 0.568 1.274 0.541 1.122 0.536
336 1.066 0.495 1.069 0.495 1.072 0.496 1.430 0.610 1.410 0.581 1.246 0.575 1.293 0.625 1.153 0.530 1.173 0.571 1.285 0.584 1.146 0.545
720 1.071 0.496 1.073 0.497 1.075 0.497 1.438 0.615 1.395 0.575 1.236 0.584 1.480 0.710 1.243 0.599 1.183 0.568 1.347 0.626 1.158 0.545
Avg. 1.057 0.494 1.060 0.496 1.060 0.496 1.414 0.606 1.387 0.562 1.170 0.550 1.378 0.662 1.155 0.542 1.166 0.570 1.295 0.569 1.134 0.541